OpenStudy (ltrout):

Write an equation of a line that goes through point (4, -5) and is parallel to line y=2x+8. SHOW ALL WORK!

OpenStudy (ltrout):

Could someone check my work for me? I did this: y=2x+8, so the slope is 2, so I did y=2x+b and filled in the numbers to get -5=2(4)+b and ended up getting b=-13 so the new equation that is parallel to y=2x+8 would be y=2x-13? Is this correct?